Consider a process consisting of three activities required to make a cell phone case. Assume there exists an unlimited demand for the cases. • Activity 1 has a processing time of 6 minutes per unit. • Activity 2 has a processing time of 3 minutes per unit. Activity 3 has a processing time of 5 minutes per unit. All three activities are staffed by one worker and each worker gets paid $12 per hour. A. Calculate the cost of direct labor? What is the meaning of direct labor to a firm?

B. Calculate the labor content? What is the labor content?

C. Calculate the average labor utilization? What is the meaning of this measure to the form?

D. Assume the demand rate is 20 units per hour-What is the takt time? What is the meaning of the takt time?
Answer #1

Answer:


a) Direct labor: It is defined as the work performed by the workers who provide the service directly to the customers, such as auditors, lawyers, and consultants.

Processing time of resource 1 is the longest.

Therefore, resource 1 is the bottleneck.

Process capacity = 60/6 = 10 units per hour

Total direct labor cost per hour = 3 workers * $ 12 per hour = $ 36 per hour

Cost of direct labor = Total direct labor cost per hour / Process capacity

= 36 / 10 = $ 3.6 per unit


b) The total labor content is defined as the time sum of all process steps.

Labor content = 6+3+5 = 14 minutes per unit


c) The average labor utilization is defined as the total labor content divided by the sum of labor content and total idle time.

Cycle time = Processing time of bottleneck resource = 6 minutes per hour

Average labor utilization = Labor cost / (Number of workers * Cycle time)

= 14 / (3*6) = 0.777 = 77.77 %


d) Takt time is the amount of time an item or service needs to be completed if a customer’s on-time delivery deadline will be met. The amount of time between one part being completed and the next part being completed must be the same or less than the takt time, or we can assume that the parts will not be produced per the customer’s schedule.

Takt Time = Customer Demand / Available Time

Demand rate = 20 units per hour

takt time = 60 minutes per hour / Demand rate

= 60/20 = 3 minutes per unit


answered by: Subrahmanyam golla